Summary Title: The search continues
Author’s name: Andrew Wheeler
Mission Status: Nominal
Sol Activity Summary: A repeat of yesterday, Sol 6 dawned bright, clear and cold; perfect weather for EVAs. It also allowed the solar observatory to be utilised for the first time. The morning EVA saw a return to Kissing Camel Ridge in search of access to Hab ridge and the locations of fossil beds, concretions and petrified wood. The return to the parked rovers during this EVA was via a different route that pushed the EVA into the reserve time. A lesson learned. The afternoon EVA chartered new territory into Candor Chasma (via Galileo Road) in the search for fossil beds and spectral targets. A location for fossil beds was identified as well as more green and red mudstone lithologies. Interestingly, purple mudstone beds and salt coated rocks were located and will be potential targets for spectral analysis on later EVAs. Pizza and fresh salad greeted all for the evening meal and very welcome it was.
Look Ahead Plan: A day of rest and consolidation. However, in expectation of clear weather again, the solar observatory will continue to have the idiosyncrasies worked through to allow clear observations of the solar surface. With no EVAs scheduled, the crew will be able to catch up on interpretations and conclusions as well as personal time.
Anomalies in work: An incursion by curious visitors in to the campus In the afternoon.
Weather: Clear crisp morning at -6 Celcius rising to low teens during the day. No precipitation. Humidity continuously dropping from near 60% to below 40%. The wind was faint to gentle during the day.
Crew Physical Status: nominal
EVA: EVA 10 with Clare and Scott to Kissing Camel Ridge en route to Hab Ridge in search of fossil beds, concretions and petrified wood.
EVA 11 to Candor Chasma via Galileo Road reconnoitering for spectral targets, fossil beds and concretions.
Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA report, EVA Request(s), Astronomer report and mid-Mission report.
Support Requested: Nothing to request.